OpenStudy (anonymous):

find the value of f(3) and f(-2) for each of the equations below: i)f(x)=(x square - 2)(x+4) f(x)=(xt3)square over (4-x)

OpenStudy (zzr0ck3r):

\((x^2-2)(x+4)\)?

OpenStudy (zzr0ck3r):

just put a 3 where you see an x \(((3)^2-2)((3)+4)\)=? the parenthesis around the 3 may seem redundant but they will help you on the next one put a-2 where you see an x \(((-2)^2-2)((-2)+4)\)=?

OpenStudy (anonymous):

\[i)f(x)=(x^{2}-2)(x+4)\]

OpenStudy (anonymous):

\[f(x)=\frac{ (x+3)^{2} }{(4-x) }\]

OpenStudy (anonymous):

answers please.